Method of Elimination solving Problem
15. Bhargav stops in at a deli to get lunch for his crew. He buys five roast beef and three vegetarian sandwiches and the order costs $27.50. The next week, he pays he pays $23.00 for two broast beef and six vegetarian sandwiches. How much does one roast beef sandwiches cost?
Please and thank you very much...

Answer by checkley77(12844)

5B+3V=27.50 MULTIPLY BY -2 & ADD.
2B+6V=23.00
-10B-6V=-55.00
-----------------
-8B=-32.00
B=-32.00/-8
B=$4.00 FOR EACH ROAST BEEF SANDWICH.
2*4.00+6V=23.00
8.00+6V=23.00
6V=23.00-8.00
6V=15.00
V=15.00/6
V=$2.50 FOR EACH VEGETARIAN SANDWICH.
PROOF:
5*4+3*2.50=27.50
20+7.50=27.50
27.50=27.50